chenoawhite1:

Proof 1: Given two intersecting lines AB and CD at point E, prove that ∠AEC≅∠DEB. (You may NOT use the Vertical Angle Theorem.) Statement Reason

chenoawhite1:



chenoawhite1:



mdobbs6856:

@Elsa213

jhonyy9:

so i think it in this way angle AEC with angle CEB equal 180 degrees so let note angle AEC an E1 and angle CEB and E2 in a second way angle AED note E4 with angle DEB note E3 equal 180 degrees but in an other way we can write that E1 +E4 = 180 degrees and that E2 +E3 = 180 degrees too so there are E1 + E2 = 180° E3 +E4 = 180° and E1 + E4 = 180° E2 + E3 = 180° how you think what result from these above wrote ? E2 need be equal E4 ? and E1 need be equal E3 ? what is your opinion about these ? hope helped understandably easy